# SLRQA: A Sparse Low-Rank Quaternion Model for Color Image Processing with Convergence Analysis

### v1

@@ 6 Numerical experiment
In this section, color image denoising and image inpainting problems are respectively used to demonstrate the performance of SLRQA in (2) with Algorithm 1 and SLRQA-NF in (4) with Algorithm 3. In Section 6.1, we give the parameter setting for SLRQA and SLRQA-NF. In Section 6.2, we compare SLRQA and SLRQA-NF models based on quaternion and RGB representation to demonstrate the rationality of quaternion representation. Furthermore, in Section 6.3 and 6.4, comparisons of SLRQA and SLRQA-NF with other state-of-the-art methods are also presented to respectively show the superiority of SLRQA and SLRQA-NF. All the experiments are performed using MATLAB R2022b running on a desktop with an Intel Core R7-5800H CPU (3.90 GHz) and 16 GB of RAM. The codes are available at https://github.com/dengzhanwang/SLRQA/tree/main.
@@ 6.1 Problem description, parameter setting, and testing environment
$\min_{{\mathbf{\dot{X}}},\mathbf{\dot{W}}\in\mathbb{H}^{m\times n}}$ $\sum_{i}\phi(\sigma_{i}({\mathbf{\dot{X}}}),\gamma)+\lambda p(\mathbf{\dot{W}})+\frac{1}{2\tau^{2}}\|{\mathbf{\dot{X}}}-\mathbf{\dot{Y}}\|_{\mathrm{F}}^{2},\quad\mathrm{s.t.}\quad\mathcal{W}({\mathbf{\dot{X}}})=\mathbf{\dot{W}},$ (55)
@@
$\min_{\mathbf{\dot{Z}},\mathbf{\dot{W}}\in\mathbb{H}^{m\times n}}$ $\sum_{i}\phi(\sigma_{i}(\mathbf{\dot{Z}}),\gamma)+\lambda p(\mathbf{\dot{W}}),\quad\mathrm{s.t.}\quad\mathcal{P}_{\Omega}(\mathcal{W}^{\#}(\mathbf{\dot{W}}))=\mathcal{P}_{\Omega}(\mathbf{\dot{Y}}),\quad\mathbf{\dot{Z}}=\mathbf{\dot{W}},$ (56)
@@
where $\mathcal{W}$ is the QDCT [12], $p(\mathbf{\dot{W}})$ is chosen as the Huber function, i.e., $(p(\mathbf{\dot{W}}))_{ij}=\frac{1}{2\delta}|\mathbf{\dot{W}}_{ij}|^{2}$ if $|\mathbf{\dot{W}}_{ij}|<\delta$; and $(p(\mathbf{\dot{W}}))_{ij}=|\mathbf{\dot{W}}_{ij}|-\frac{1}{2}\delta$ otherwise, $\delta>0$ is a prescribed constant, and the function $\phi$ are chosen as the Schatten-$\gamma$, Laplace, or Weighted Schatten-$\gamma$ functions. Note that the color image denoising problem (55) is the SLRQA in (2) with $\mathcal{A}$ being identity, $\Omega$ being the entire indices of pixels, and $p$ being the Huber function. The color image inpainting problem (56) is the SLRQA-NF in (4) with $\mathcal{A}$ being identity and $p$ being the Huber function. Problem (55) with the Schatten-$\gamma$, Laplace, and Weighted Schatten-$\gamma$ functions for $\phi$ is respectively denoted by SLRQA-1, SLRQA-2, and SLRQA-3. Likewise, Problem (56) using the three options of $\phi$ is denoted by SLRQA-NF-1, SLRQA-NF-2, and SLRQA-NF-3.
@@
The ten images in Figure 1 are used to generate simulation problems. Specifically, for image denoising, the noisy images $\mathbf{\dot{Y}}$ are obtained by adding additive white Gaussian noise with zero mean and variance $\tau^{2}$ to the color images in Figure 1, and for image inpainting, each pixel of an image is chosen to be in the indices set $\Omega$ with the same probability $\chi\in(0,1)$. Therefore, $\chi$ is also called the missing rate. The parameters $\tau$, $\lambda$, and $\gamma$ are specified later when reporting numerical results. Note that the parameters $\gamma$ and $\lambda$ depend on the choice of the surrogate function $\phi$, the noise level, and whether the non-local self-similarity is used. The parameter $\delta$ is set depending on the accuracy of the current iterate. Specifically, $\delta=1$ if $\epsilon_{k}\geq 10^{-2}$, $\delta=10^{-2}$ if $10^{-3}\leq\epsilon_{k}<10^{-2}$, and $\delta=10^{-4}$ otherwise, where $\epsilon_{k}=\|{\mathbf{\dot{X}}}_{k+1}-{\mathbf{\dot{X}}}_{k}\|_{\mathrm{F}}+\|\mathbf{\dot{W}}_{k+1}-\mathbf{\dot{W}}_{k}\|_{\mathrm{F}}+\|\mathcal{W}({\mathbf{\dot{X}}}_{k+1})-\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}$ in Algorithm 1 and $\epsilon_{k}=\|\mathbf{\dot{Z}}_{k+1}-\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}+\|\mathcal{P}_{\Omega}(\mathcal{W}^{\#}(\mathbf{\dot{W}}_{k+1}))-\mathcal{P}_{\Omega}(\mathbf{\dot{Y}})\|_{\mathrm{F}}$ in Algorithm 3.
@@
The parameters in Algorithm 1, 2 and 3 are set as $\mu=1.1$, $\beta=10$, $L_{1,k}=L_{2,k}=1,\forall k$, $\eta_{C}=10^{-10}$, and $\eta=10^{-4}$. Note that $\eta_{C}=10^{-10}$ is for the high accuracy of the ${\mathbf{\dot{X}}}$ or $\mathbf{\dot{Z}}$-subproblem.
@@
Peak Signal-to-Noise Ration (PSNR) and Structure Similarity (SSIM) [30] are chosen to measure the quality of the images recovered by all methods. Specifically, the PSNR of two color image $\mathbf{X}$ and $\mathbf{Y}\in\mathbb{R}^{m\times n\times 3}$ is calculated as: ${\rm PSNR}(\mathbf{X},\mathbf{Y})=10\log_{10}\frac{\textit{peakval}}{\rm{MSE}},$ where $\text{MSE}=\frac{1}{3mn}\sum_{i=1}^{m}\sum_{j=1}^{n}\sum_{k=1}^{3}|\mathbf{X}_{i,j,k}-\mathbf{Y}_{i,j,k}|^{2},$ and peakval is taken from the range of the image data type. Furthermore, the SSIM of $\mathbf{X}$ and $\mathbf{Y}$ is given by
$\text{SSIM}(\mathbf{X}_{ij},\mathbf{Y}_{ij})=\frac{1}{MN}\sum_{i=1}^{M}\sum_{j=1}^{N}\text{SSIM}(\mathbf{X}_{ij},\mathbf{Y}_{ij})$ (57)
$\text{SSIM}(\mathbf{X}_{ij},\mathbf{Y}_{ij})=l(\mathbf{X}_{ij},\mathbf{Y}_{ij})\,c(\mathbf{X}_{ij},\mathbf{Y}_{ij})\,s(\mathbf{X}_{ij},\mathbf{Y}_{ij}),$ (58)
where $\mathbf{X}_{ij}$ and $\mathbf{Y}_{ij}$ denote the local image patches, $M$ and $N$ denote the number of local patches along horizontal and vertical direction respectively, the three functions $l(\mathbf{X}_{ij},\mathbf{Y}_{ij})$, $c(\mathbf{X}_{ij},\mathbf{Y}_{ij})$, and $s(\mathbf{X}_{ij},\mathbf{Y}_{ij})$ respectively denote luminance item, contrast item, and structure item, and they are defined as follows:
$l(\mathbf{X}_{ij},\mathbf{Y}_{ij})=\frac{2\mu_{\mathbf{X}_{ij}}\mu_{\mathbf{Y}_{ij}}+C_{1}}{\mu_{\mathbf{X}_{ij}}^{2}+\mu_{\mathbf{Y}_{ij}}^{2}+C_{1}},\quad c(\mathbf{X}_{ij},\mathbf{Y}_{ij})=\frac{2\sigma_{\mathbf{X}_{ij}}\sigma_{\mathbf{Y}_{ij}}+C_{2}}{\sigma_{\mathbf{X}_{ij}}^{2}+\sigma_{\mathbf{Y}_{ij}}^{2}+C_{2}},\quad s(\mathbf{X}_{ij},\mathbf{Y}_{ij})=\frac{\sigma_{\mathbf{X}_{ij}\mathbf{Y}_{ij}}+C_{3}}{\beta_{\mathbf{X}_{ij}}\sigma_{\mathbf{Y}_{ij}}+C_{3}},$ (59)
with $\mu_{\mathbf{X}_{ij}},\mu_{\mathbf{Y}_{ij}},\sigma_{\mathbf{X}_{ij}},\sigma_{\mathbf{Y}_{ij}},\sigma_{{\mathbf{X}_{ij}}{\mathbf{Y}_{ij}}},C_{1},C_{2}$ and $C_{3}$ being the mean, standard deviation and cross-covariance of image patches $\mathbf{X}_{ij}$ and $\mathbf{Y}_{ij}$, $10^{-4},9\times 10^{-4}$ and $4.5\times 10^{-4}$ respectively. MATLAB commands psnr and ssim are used to compute the PSNR and SSIM for color images. In general, the higher values the PSNR and SSIM are, the better the denoising quality is.
@@ 6.2.1 Image denoising
The PSNR and SSIM values of SLRQA with quaternion and RGB representation are reported in Table 2. Furthermore, the comparison of the two representations for SLRQA-1 with different noise levels is shown in Figure 2. The figure and table show that SLRQA with quaternion representation achieves higher PSNR and SSIM compared with RGB representation in all cases. This observation coincides with the intuition that quaternion representation utilizes the information between channels to achieve better performance.
@@ 6.3 Color image denoising
NSS is a significant technique in color image denoising problems. The comparison without NSS directly reflects the superiority of the compared models, and the comparison with NSS better reflects the robustness of the models in practice. To fully compare the SLRQA with other state-of-the-art color image denoising methods, we tested their performance without and with the usage of non-local self-similarity (NSS) respectively in Section 6.3.1 and Section 6.3.2.
@@ 6.3.1 Color image denoising without NSS
The compared methods include SRRC[39], WNNM[14], LRQA[10], and DWT[5]333 The implementations that we use are available from:
@@
WNNM: http://www4.comp.polyu.edu.hk/ cslzhang/papers.htm
@@
. SRRC exploits the weighted Schatten $p$-norm as the regularizer for the rank residual to obtain a new rank minimization model. WNNM uses the weighted nuclear norm as an alternative function to the rank function. LRQA is a quaternion-based model that makes use of the low-rankness property by a nonconvex surrogate function. Compared with SLRQA, none of SRRC, WNNM, and LRQA exploits the sparsity feature of the image. DWT [5] is a vanilla analysis-based approach (5) where low-rankness prior is not used and is implemented on our own. It is noted that SRRC, WNNM, and DWT are RGB-based methods, and LRQA and SLRQA are quaternion-based models. LRQA-1 and LRQA-2 denote the LRQA models using Schatten-$\gamma$ and Laplace function, respectively. For the compared methods, we use the default parameters stated in the corresponding papers.
The PSNR and SSIM results of the compared methods are shown in Table 4. It can be seen that SLRQA has better results than the other methods in terms of PSNR and SSIM in most cases. Furthermore, The PSNR value of SRRC, WNNM, and DWT are much lower than that of SLRQA when $\tau$ is large. Note that since NSS is not utilized in these methods, the results in Table 4 underscore the effectiveness and robustness of the SLRQA model and further validate the advanced nature and adaptability of its intrinsic architecture. We also give the visual comparison between all competing methods and SLRQA as shown in Figure 4. It can be seen in the highlighted red rectangles that when compared with other denoising methods, SLRQA can preserve more details of images. In particular, in terms of hair details, the images denoised by the DWT and LRQA-1 methods are still noisy. The images recovered by LRQA and SRRC methods contain blurry parts and hence the detail features of the images are lost. In contrast, the images recovered by SLRQA models have better results both numerically and visually.
To further analyze the details of PL-ADMM, we plot the empirical convergence of PL-ADMM for Problem (55) on ten tested images with $\tau=30$. The results are shown in Figure 5. It can be observed that the error curve has a downward trend, illustrating the empirical convergence of Algorithm 1. We can observe that $\|\Delta{\mathbf{\dot{X}}}_{k+1}\|_{\mathrm{F}}+\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}+\|\Delta\mathbf{\dot{\Lambda}}_{k+1}\|_{\mathrm{F}}\rightarrow 0$. Hence, it follows from (42) that $|\|\widetilde{\mathbf{\dot{D}}}_{k+1}\||\rightarrow 0$. According to the update of $\mathbf{\dot{\Lambda}}_{k+1}$ and Lemma 4.7, it follows that $({\mathbf{\dot{X}}},\mathbf{\dot{W}},\mathbf{\dot{\Lambda}})$ converges to a stationary point of $\mathcal{L}_{\beta}$ and hence a KKT point of the original problem in (55).

### v2

@@ 5 Numerical experiment
In this section, color image denoising and image inpainting problems are respectively used to demonstrate the performance of SLRQA in (2) with Algorithm 1 and SLRQA-NF in (4) with Algorithm 3. In Section 5.1, we give the parameter setting for SLRQA and SLRQA-NF. In Section 5.2, we compare SLRQA and SLRQA-NF based on quaternion and RGB representation to demonstrate the rationality of quaternion representation. Furthermore, in Section 5.3 and 5.4, comparisons of SLRQA and SLRQA-NF with other state-of-the-art methods are also presented to respectively show the superiority of SLRQA and SLRQA-NF. All the experiments are performed using MATLAB R2022b running on a desktop with an Intel Core R7-5800H CPU (3.90 GHz) and 16 GB of RAM. The codes are available at https://github.com/dengzhanwang/SLRQA/tree/main.
@@ 5.1 Problem description, parameter setting, and testing environment
$\min_{{\mathbf{\dot{X}}},\mathbf{\dot{W}}\in\mathbb{H}^{m\times n}}$ $\sum_{i}\phi(\sigma_{i}(\sqrt{{{\mathbf{\dot{X}}}}^{*}{\mathbf{\dot{X}}}+\varepsilon^{2}\mathbf{I}}),\gamma)+\lambda p(\mathbf{\dot{W}})+\frac{1}{2\tau^{2}}\|{\mathbf{\dot{X}}}-\mathbf{\dot{Y}}\|_{\mathrm{F}}^{2},\quad\mathrm{s.t.}\quad\mathcal{W}({\mathbf{\dot{X}}})=\mathbf{\dot{W}},$ (88)
@@
$\min_{\mathbf{\dot{Z}},\mathbf{\dot{W}}\in\mathbb{H}^{m\times n}}$ $\sum_{i}(\sigma_{i}(\sqrt{{\mathbf{\dot{Z}}}^{*}\mathbf{\dot{Z}}+\varepsilon^{2}\mathbf{I}}),\gamma)+\lambda p(\mathbf{\dot{W}}),\quad\mathrm{s.t.}\quad\mathcal{P}_{\Omega}(\mathcal{W}^{\#}(\mathbf{\dot{W}}))=\mathcal{P}_{\Omega}(\mathbf{\dot{Y}}),\quad\mathbf{\dot{Z}}=\mathbf{\dot{W}},$ (89)
@@
where $\mathcal{W}$ is the QDCT, see details in [14], $p(\mathbf{\dot{W}})$ is chosen as the Huber function, and the function $\phi$ is chosen as the Schatten-$\gamma$, Laplace, or Weighted Schatten-$\gamma$ functions. Note that the color image denoising problem (88) is SLRQA in (2) with $\mathcal{A}$ being identity, $\Omega$ being the entire indices of pixels, and $p$ being the Huber function. The color image inpainting problem (89) is SLRQA-NF in (4) with $\mathcal{A}$ being identity and $p$ being the Huber function. Problem (88) with the Schatten-$\gamma$, Laplace, and Weighted Schatten-$\gamma$ functions for $\phi$ is respectively denoted by SLRQA-1, SLRQA-2, and SLRQA-3. Likewise, Problem (89) using the three options of $\phi$ is denoted by SLRQA-NF-1, SLRQA-NF-2, and SLRQA-NF-3.
@@
The ten images in Figure 1 are used to generate simulation problems. Specifically, for image denoising, the noisy images $\mathbf{\dot{Y}}$ are obtained by adding additive white Gaussian noise with zero mean and variance $\tau^{2}$ to the color images in Figure 1, and for image inpainting, each pixel of an image is chosen to be in the indices set $\Omega$ with the same probability $\chi\in(0,1)$. Therefore, $\chi$ is also called the missing rate. The parameters $\tau$, $\lambda$, and $\gamma$ are specified later when reporting numerical results. Note that the parameters $\gamma$ and $\lambda$ depend on the choice of the surrogate function $\phi$, the noise level, and whether the non-local self-similarity is used. The parameter $\delta$ is set depending on the accuracy of the current iterate. Specifically, $\delta=1$ if $\epsilon_{k}\geq 10^{-2}$, $\delta=10^{-2}$ if $10^{-3}\leq\epsilon_{k}<10^{-2}$, and $\delta=10^{-4}$ otherwise, where $\epsilon_{k}=\|{\mathbf{\dot{X}}}_{k+1}-{\mathbf{\dot{X}}}_{k}\|_{\mathrm{F}}+\|\mathbf{\dot{W}}_{k+1}-\mathbf{\dot{W}}_{k}\|_{\mathrm{F}}+\|\mathcal{W}({\mathbf{\dot{X}}}_{k+1})-\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}$ in Algorithm 1 and $\epsilon_{k}=\|\mathbf{\dot{Z}}_{k+1}-\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}+\|\mathcal{P}_{\Omega}(\mathcal{W}^{\#}(\mathbf{\dot{W}}_{k+1}))-\mathcal{P}_{\Omega}(\mathbf{\dot{Y}})\|_{\mathrm{F}}$ in Algorithm 3. Furthermore, for PL-ADMM-NF, the objective function of the subproblem of $\mathbf{\dot{W}}$ is differentiable, we solve the subproblem using the gradient descent method.
@@
The parameters in Algorithm 1, 2 and 3 are set as $\mu=1.1$, $\beta=10$, $L_{1,k}=L_{2,k}=1,\forall k$, $\eta_{C}=10^{-10}$, and $\eta=10^{-4}$. Note that $\eta_{C}=10^{-10}$ is for the high accuracy of the ${\mathbf{\dot{X}}}$ or $\mathbf{\dot{Z}}$-subproblem. The threshold $\varepsilon$ of $\phi$ is set to $10^{-2}$ in our experiments.
@@
Peak Signal-to-Noise Ration (PSNR) and Structure Similarity (SSIM) [36] are chosen to measure the quality of the images recovered by all methods. MATLAB commands psnr and ssim are used to compute the PSNR and SSIM for color images. In general, the higher values the PSNR and SSIM are, the better the denoising quality is.
@@ 5.2.1 Image denoising
The PSNR and SSIM values of SLRQA with quaternion and RGB representation are reported in Table 2. Furthermore, the comparison of the two representations for SLRQA-1 with different noise levels is shown in Figure 2. The figure and table show that SLRQA with quaternion representation achieves higher PSNR and SSIM compared with RGB representation in all cases. This observation coincides with the intuition that quaternion representation uses the information between channels to achieve better performance.
@@ 5.3 Color image denoising
Non-local self-similarity (NSS) is a significant technique in color image denoising problems. The comparison without NSS directly reflects the superiority of the compared models, and the comparison with NSS better reflects the robustness of the models in practice. To fully compare the SLRQA with other state-of-the-art color image denoising methods, we tested their performance without and with the usage of NSS respectively in Section 5.3.1 and Section 5.3.2.
@@ 5.3.1 Color image denoising without NSS
The compared methods include pQSTV [37], QWSNM [48], MCNNFNM [33], SRRC [49], QWNNM [20], LRQA [12], and DWT [7]333 The implementations that we use are available at
pQSTV: https://github.com/Huang-chao-yan/pQSTV.git
QWSNM: https://github.com/qiuxuanzhizi/QWSNM.git
MCNNFNM: https://github.com/wangzhi-swu/MCNNFNM.git
@@
. QWSNM, MCNNFNM, SRRC, QWNNM, and LRQA are low-rank minimization models. QWSNM extends the WSNM into the quaternion domain for color restoration. MCNNFNM is a multi-channel nuclear norm minus Frobenius norm minimization model. SRRC exploits the weighted Schatten $p$-norm as the regularizer for the rank residual to obtain a new rank minimization model. QWNNM uses the weighted nuclear norm as an alternative function to the rank function. LRQA is a quaternion-based model that makes use of the low-rankness property by a nonconvex surrogate function. Compared with SLRQA, none of SRRC, QWNNM, and LRQA exploit the sparsity feature of the image. pQSTV is a dictionary method for color image denoising that combines total variation regularization and pure quaternion representation. DWT [7] is a vanilla analysis-based approach (5) where low-rankness prior is not used and is implemented on our own. It is noted that MCNNFNM, SRRC, and DWT are monochromatic methods, while QWNNM, pQSTV, QWSNM, LRQA, and SLRQA are quaternion-based models. LRQA-1 and LRQA-2 denote the LRQA models using Schatten-$\gamma$ and Laplace function, respectively. The comparisons with LRQA and QWNNM can be regarded as ablation experiments to present the significance of sparsity and low-rankness prior respectively. For the compared methods, we use the default parameters stated in the corresponding papers.
The PSNR and SSIM results of the compared methods are shown in Table 4. It can be seen that SLRQA has better results than the other methods in terms of PSNR and SSIM in most cases. Furthermore, the PSNR values of SRRC, QWNNM, and DWT are much lower than that of SLRQA when $\tau$ is large. Note that since NSS is not used in these methods, the results in Table 4 underscore the effectiveness and robustness of the SLRQA and further validate the advanced nature and adaptability of its intrinsic architecture. We also give the visual comparison between all competing methods and SLRQA as shown in Figure 4. It can be seen in the highlighted red rectangles that when compared with other denoising methods, SLRQA can preserve more details of images. In particular, in terms of hair details, the images denoised by the DWT and LRQA-1 methods are still noisy. The images recovered by LRQA and SRRC methods contain blurry parts, and hence the detailed features of the images are lost. In contrast, the images recovered by SLRQAs have better results both numerically and visually.
To further analyze the details of PL-ADMM, we plot the empirical convergence of PL-ADMM for Problem (88) on ten tested images with $\tau=30$. The results are shown in Figure 5. It can be observed that the error curve has a downward trend, illustrating the empirical convergence of Algorithm 1. We can observe that $\|\Delta{\mathbf{\dot{X}}}_{k+1}\|_{\mathrm{F}}+\|\Delta\mathbf{\dot{W}}_{k+1}\|_{\mathrm{F}}+\|\Delta\mathbf{\dot{\Lambda}}_{k+1}\|_{\mathrm{F}}\rightarrow 0$. Hence, it follows that $({\mathbf{\dot{X}}},\mathbf{\dot{W}},\mathbf{\dot{\Lambda}})$ converges to a stationary point of $\mathcal{L}_{\beta}$ and hence a KKT point of the original problem in (88) [43, Theorem 2].